OpenStudy (anonymous):

is there a difference between f(x)=log(x)-2 and f(x)=log(x-2) on the graph?

OpenStudy (anonymous):

the issue is that i have ran both on a calculator and come up with two different (yet similar) graphs....but of course both is offered as answers

OpenStudy (anonymous):

so here is f(x)=log(x-2)

OpenStudy (anonymous):

here is f(x)=log(x)-2

OpenStudy (anonymous):

perfect, that is what i found...with y=f(x), it moves ot the right...which would be the first graph...ty

OpenStudy (anonymous):

your welcome :)